The shortest wave length emitted by bat about 3.3mm.what is corresponding frequency


1
Expert's answer
2021-02-08T18:37:58-0500

We can find the frequency from the wave speed equation:


"v=f\lambda,""f=\dfrac{v}{\lambda}=\dfrac{340\ \dfrac{m}{s}}{3.3\cdot10^{-3}}=103\ kHz."